OpenStudy (anonymous):

What is the solution of the following system? x - y =11 -x + y = -11

OpenStudy (ivettef365):

what happen when you subtract both equation x -y = 11 -x +y = -11 ----------- ?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Does this result mean there is no solution?

OpenStudy (anonymous):

When you graph it out, they are the same line. Because of this, there is no intersection of the lines which results in no solution

OpenStudy (ivettef365):

in the other hand it have many solutions

OpenStudy (austinl):

Infinitely many solutions.

OpenStudy (anonymous):

thanks guys ! i forgot that and my precalc teacher just reminded us this week haha

OpenStudy (anonymous):

Thank you everyone for your help!

OpenStudy (ivettef365):

:)